Adam is half as old as Erin. The difference in their ages is 16. Erin's age is the variable x.
Answer by stanbon(75887)

Adam is half as old as Erin. The difference in their ages is 16. Erin's age is the variable x.
--------------------------------
Equations:
A = (1/2)E
E - A = 16
----------------
Substitute for "A" and solve for E:
E - (1/2)E = 16
(1/2)E = 16
E = 32 (Erin's age)
A = (1/2)32 = 16 (Adam's age)
=================================
Cheers,
Stan H.